You are filling a cylindrical drum (D = 22.5 in., H = 33.5 in.) with water flowing from a garden faucet at a rate of 10 GPM. How long will you have to wait to fill the drum? Express you answer in minutes.




1
Expert's answer
2021-11-29T11:40:53-0500

First, remember that 1 GPM (1 gallon per minute) is 231 in3/min.

Find the volume:


"V=AH=\frac{\pi D^2}4H=13320\text{ in}^3."

The time:


"t=\frac Vq=\frac{13320}{10·231}=5.77\text{ min}."